OpenStudy (anonymous):

Which proportion best represents the following? John bought apples at a rate of 0.5 lbs for $3.50. At this rate, how much would 2.8 lbs cost

OpenStudy (anonymous):

can somebodys pls help

OpenStudy (anonymous):

I would use a ratio: 0.5 lbs : $3.50 0.5:3.5 2.8:19.6 (multiply both sides by 5.6) Therefore it would cost $19.6 for 2.8lbs

OpenStudy (anonymous):

awnsers are 0.5/3.50=x/2,8 0.5/2.8=x/3.50 0.5/3.50=2.8/x

OpenStudy (anonymous):

Your answers are formatted as : pound : cost Cost is your x, therfore your answer must be in the format: pound/x

OpenStudy (anonymous):

Those are just the two ratios expressed as an equation. We have 0.5:3.5 and 2.8:x (because 19.6 is unknown until you do calculation) In order to obtain the answer using algebra you can write the two ratios like this: \[\frac{ 0.5 }{ 3.5 }=\frac{ 2.8 }{ x }\]

OpenStudy (anonymous):

thank you riddellikins

OpenStudy (anonymous):

You are very welcome